OpenStudy (anonymous):

how many adults must be randomly selected to estimate the mean FICO score of working adults in the United States. We want 95% confidence that the sample mean is within 3 points of the population mean, and the population standard deviation is 68

OpenStudy (anonymous):

What formula do you know? Is this something like Chebyshev?

OpenStudy (kropot72):

Substitute the given values into the following formula and solve for n (the sample size). \[\bar {x}-1.96\frac{\sigma}{\sqrt{n}}<\mu < \bar {x}+1.96\frac{\sigma}{\sqrt{n}}\]